# PF-LHM: 3D Animatable Avatar Reconstruction from Pose-free Articulated Human Images

### v1

@@ Abstract
Reconstructing an animatable 3D human from casually captured images of an articulated subject without camera or human pose information is a practical yet challenging task due to view misalignment, occlusions, and the absence of structural priors. While optimization-based methods can produce high-fidelity results from monocular or multi-view videos, they require accurate pose estimation and slow iterative optimization, limiting scalability in unconstrained scenarios. Recent feed-forward approaches enable efficient single-image reconstruction but struggle to effectively leverage multiple input images to reduce ambiguity and improve reconstruction accuracy. To address these challenges, we propose PF-LHM, a large human reconstruction model that generates high-quality 3D avatars in seconds from one or multiple casually captured pose-free images. Our approach introduces an efficient Encoder-Decoder Point-Image Transformer architecture, which fuses hierarchical geometric point features and multi-view image features through multimodal attention. The fused features are decoded to recover detailed geometry and appearance, represented using 3D Gaussian splats. Extensive experiments on both real and synthetic datasets demonstrate that our method unifies single- and multi-image 3D human reconstruction, achieving high-fidelity and animatable 3D human avatars without requiring camera and human pose annotations. Code and models will be released to the public.
Figure 1: 3D Avatar Reconstruction and Animation Results of our PF-LHM. Given a set of $N\geq 1$ images of a human subject, without requiring camera parameters or human pose annotations, our method can reconstruct a high-fidelity, animatable 3D human avatar in seconds.
@@ 1 Introduction
In this work, we propose PF-LHM, a novel feed-forward framework for fast and high-fidelity 3D human reconstruction from one or a few images without requiring the camera and human poses. To achieve this, we design an efficient Encoder-Decoder Point-Image Transformer (PIT) Framework that hierarchically fuses 3D geometric features with multi-image cues. The framework is built upon Point-Image Transformer blocks (PIT-blocks), which enable interaction between geometric and image tokens via attention fusion while maintaining scalability through spatial hierarchy.
@@
The resulting 3D geometry tokens are decoded to predict Gaussian splatting parameters, enabling photorealistic rendering and animation. To enhance robustness and generalization, we train our model on large-scale real-world human video datasets covering diverse clothing styles, body shapes, and viewing conditions.
In summary, our contributions are:
  • We introduce PF-LHM, to the best of our knowledge, the first feed-forward model capable of reconstructing high-quality, animatable 3D human avatars in seconds from one or a few casually captured images, without requiring either camera poses or human pose annotations.
@@
  • Extensive experiments on both synthetic and real-world data demonstrate that PF-LHM unifies single- and multi-image 3D human reconstruction, with superior generalization and visual quality.
@@ 2 Related Work
2.1 Human Reconstruction from A Single Image
@@
Inspired by the success of large reconstruction models [10, 37], emerging solutions aim to enable direct generalizable reconstruction through feed-forward networks which significantly accelerate the inference time. Human-LRM [46] employs a feed-forward model to decode the triplane NeRF representation, then followed by a conditional diffusion-based novel views generation and reconstruction. IDOL [60] introduces a UV-Alignment transformer model to decode Gaussian attribute maps in a structured 2D UV space. LHM [31] leverages a Body-Head multimodal transformer architecture produces animatable 3D avatars with the face identity preservation and fine detail recovery. While these single-view methods often face challenges with occlusions and invisible regions, frequently resulting in geometrically implausible results or blurred textures, the proposed PF-LHM leverages a variable number of pose-free images to reconstruct photorealistic and animatable avatars.
Concurrent work, GIGA [61], introduces a generalizable human reconstruction model based on UV map representations. However, in contrast to our approach, GIGA requires sparse-view inputs where the same action is captured from multiple viewpoints, along with complex camera setups and motion calibration. These constraints make it challenging to apply in causal scenarios.
2.2 Human Reconstruction from Monocular Videos
@@
Another series of works [14, 23, 11, 8, 15, 29, 55] take it further by incorporating a 3D parametric human model into the optimization process, and thus can drive the human reconstruction in novel poses without any post-processing. Despite impressive visual fidelity, they often require dozens of minutes and dozens of views for a good optimization, which limits their practical usage in real-world scenarios.
Unconstrained collection is ideal input for a practical application. However, existing methods [51, 54] share a similar pipeline that uses a view generative model and score distillation sampling [27] for shape optimization. As a result, they are costly for offline training and impractical for online reconstruction. Paving in a new way, PF-LHM infers a human avatar from one, a few, or dozens of views under any poses in a feed-forward manner and costs only seconds, which is extremely efficient for online applications. Moreover, PF-LHM greatly outperforms any previous state-of-the-art methods on 3D human reconstruction and offers a more flexible input manner for the community.
2.3 Feed-Forward Scene Reconstruction
Recent years have witnessed a paradigm shift in geometric 3D vision, driven by the emergence of methods that eliminate traditional dependencies on camera calibration and multi-stage pipelines. At the forefront of this revolution lies the DUSt3R [42] framework, which reimagines 3D reconstruction as a direct regression problem from image pairs to 3D pointmaps. By discarding the need for intrinsic camera parameters, extrinsic pose estimation, or even known correspondence relationships, DUSt3R and its successors [52, 38, 22, 40] have democratized 3D vision, enabling rapid reconstruction across diverse scenarios while achieving state-of-the-art performance in depth estimation, relative pose recovery, and scene understanding. However, general feed-forward reconstruction methods assume that images are captured from a static scene [41, 18], while our PF-LHM can accept multiple human images with different camera and human poses as input and produce an animatable 3D avatar.
Table 1: Comparison with state-of-the-art 3D human reconstruction methods. FF stands for Feed-forward, PF for Pose-free, and AM for Animatable.
Method # Image FF PF AM Runtime
IDOL [60] $1$ Seconds
AniGS [32] $1$ 15 Minutes
LHM [31] $1$ Seconds
Vid2Avatar [7] $>100$ 1-2 Days
GPS-Gaussian [58] $2$ Seconds
3DGS-Avatar [28] $>20$ 0.5 Hours
InstantAvatar [14] $>20$ Minutes
GaussianAvatar [11] $>20$ 0.5-6 Hours
ExAvatar [23] $>20$ 1.5-5 Hours
@@ 2.3 Feed-Forward Scene Reconstruction
Vid2Avatar-Pro [8] $>100$ Hours
@@
Figure 2: Overview of the proposed PF-LHM. In the 2D space, we extract image tokens $\mathbf{T}_{\text{Img}}$ by DINOv2 from the input RGB images without camera parameters or human poses, which are then concatenated with deformation tokens $\mathbf{T}_{\text{Def}}$ to form 2D tokens $\mathbf{T}_{\text{2D}}$. In the 3D space, geometric tokens $\mathbf{T}_{\text{3D}}$ are represented by the MLP output of SMPL-X anchor points. Subsequently, we build our Encoder-Decoder Point-Image Transformer (PIT) to hierarchically fuse 3D tokens with 2D tokens, where the downsampled 3D tokens interact with 2D tokens via multimodal attention in each layer. The finalized 3D tokens are decoded to directly predict 3D Gaussian parameters, enabling animation and photorealistic rendering.
@@ 3.1 Overview
Given a set of $N\geq 1$ RGB images of a human subject, without known camera parameters or human pose annotations, our goal is to reconstruct a high-fidelity, animatable 3D human avatar in seconds.
@@
$\mathbf{\chi}\{\mathbf{p},\mathbf{r},\mathbf{f},\rho,\sigma\}=\text{PF-LHM}(P\mid I^{1},\dots,I^{N}).$ (1)
@@
To mitigate this issue, we explore strategies to reduce the number of geometric point tokens involved in attention. However, we empirically observe that simply reducing the number of point tokens significantly degrades reconstruction performance. To address this trade-off, we propose an efficient Encoder-Decoder Point-Image Transformer Framework to fuse image features with geometric point features, as illustrated in Fig. 2, which maintains reconstruction quality while reducing the attention footprint.
The final geometric point features output from the decoder are utilized to regress 3D Gaussian parameters using lightweight multi-layer perceptron (MLP) heads. To account for non-rigid deformations such as clothing or hair, we introduce learnable deformation-aware tokens. These tokens, together with the geometric features, are used to predict residual offsets in canonical space. Finally, Linear Blend Skinning (LBS) is applied to animate the canonical avatar into the target pose.
@@ 3.2 Encoder-Decoder Point-Image Transformer Framework
We begin by projecting the SMPL-X anchor points in canonical space into a set of geometric tokens and encoding the input images into image tokens, as described in Sec. 3.3. The network consists of $N_{\text{layer}}$ PIT blocks, divided into an encoder stage and a decoder stage. In the first $\lfloor N_{\text{layer}}/2\rfloor$ encoder blocks, we progressively reduce the spatial resolution of the geometric tokens using Grid Pooling [47]. At each layer, the downsampled point tokens perform the attention operation with the image tokens, enabling compact geometric representations enriched with multi-image visual cues.
@@ 3.3 Geometric Point and Images Tokenization
To incorporate human body priors, we initialize a set of 3D query points $\{\mathbf{x}_{i}\}_{i=1}^{N_{\text{points}}}\subset\mathbb{R}^{3}$ by uniformly sampling from the mesh of a canonical SMPL-X pose. Following the design of Point Transformer v3 (PTv3) [48], we first serialize these points into a structured sequence and then project them into a higher-dimensional feature space using an MLP. Formally, this process is expressed as:
@@
$\mathbf{T}_{\text{Img}}=\mathcal{E}_{\text{Img}}(I)\in\mathbb{R}^{N_{\text{I}}\times C},$ (3)
@@
Deformation-aware Token Injection
To account for non-rigid deformations such as clothing and hair present in each image, we introduce a learnable deformation token specific to the observed subject, denoted as $\mathbf{T}_{\text{Def}}\in\mathbb{R}^{1\times C}$. This token is concatenated with the image token sequence $\mathbf{T}_{\text{Img}}$, forming the multi-image tokens:
$\mathbf{T}_{\text{I}}=[\mathbf{T}_{\text{Img}};\mathbf{T}_{\text{Def}}]\in\mathbb{R}^{(N_{\text{I}}+1)\times C},$ (4)
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes token-wise concatenation along the sequence dimension.
@@ 3.4 Point-Image Transformer Block
Given the image token sequence $\mathbf{T}_{\text{2D}}=\{\mathbf{T}^{1}_{\text{I}},\dots,\mathbf{T}^{N}_{\text{I}}\}\in\mathbb{R}^{N\times(N_{\text{I}}+1)\times C}$, we apply self-attention independently to the tokens of each image. This updates the features within each frame based on its own image tokens:
$\mathbf{T}_{\text{2D}}=\text{Self-Attention}(\mathbf{T}_{\text{2D}}).$ (6)
@@
$\mathbf{\bar{T}}_{\text{2D}}$ $=\text{Flatten}(\mathbf{T}_{\text{2D}})\in\mathbb{R}^{N(N_{\text{I}}+1)\times C},$ (7)
$\mathbf{\bar{T}}_{\text{3D}}$ $=\text{MLP}_{\text{proj}}(\mathbf{T}_{\text{3D}})\in\mathbb{R}^{N_{\text{points}}\times C},$
[…]
$\mathbf{T}_{\text{3D}}$ $=\text{MLP}_{\text{uproj}}(\mathbf{\bar{T}}_{\text{3D}})\in\mathbb{R}^{N_{\text{points}}\times C_{\text{point}}},$
$\mathbf{T}_{\text{2D}}$ $=\text{UnFlatten}(\mathbf{\bar{T}}_{\text{2D}},N)\in\mathbb{R}^{N\times(N_{\text{I}}+1)\times C}.$
This global point-image attention module enables effective fusion of geometric and visual features by leveraging cross-modal attention.
@@
Pose Conditioned Deformation
Although the regressed canonical-space Gaussians can be animated to target poses using Linear Blend Skinning (LBS), modeling clothing deformations presents challenges due to their complex non-rigid motion patterns, which LBS is often unable to capture adequately. To overcome this limitation, we use a lightweight MLP to predict pose-dependent residual deformations.
Specifically, we derive a deformation-aware token $\mathbf{\bar{T}}_{\text{def}}$ by averaging the fused deformation-aware tokens $\mathbf{T}_{\text{def}}$ across all frames, and concatenate it with the SMPL parameters to modulate the geometric tokens using Adaptive Layer Normalization. These modulated features are then processed through a sequence of MLP layers to generate non-rigid residual deformations:
$\Delta\mathbf{p}_{i}^{\text{motion}}=\text{MLP}_{\text{motion}}\left(\text{AdaLN}(\mathbf{T}_{\text{points}}^{(i)},[\mathbf{\bar{T}}_{\text{def}};\bm{\theta}])\right),$ (9)
where $\bm{\theta}$ represents the SMPL pose parameters. The final posed positions are then obtained by adding both canonical offsets and motion-specific deformations before LBS is applied.

### v2

# LHM++: An Efficient Human Reconstruction Model for Pose-free Images to 3D

@@ Abstract
Reconstructing animatable 3D humans from casually captured images of articulated subjects without camera or pose information is highly practical but remains challenging due to view misalignment, occlusions, and the absence of structural priors. In this work, we present LHM++, an efficient large-scale human reconstruction model that generates high-quality, animatable 3D avatars within seconds from one or multiple pose-free images. At its core is an Encoder–Decoder Point–Image Transformer architecture that progressively encodes and decodes 3D geometric point features to improve efficiency, while fusing hierarchical 3D point features with image features through multimodal attention. The fused features are decoded into 3D Gaussian splats to recover detailed geometry and appearance. To further enhance visual fidelity, we introduce a lightweight 3D-aware neural animation renderer that refines the rendering quality of reconstructed avatars in real time. Extensive experiments show that our method produces high-fidelity, animatable 3D humans without requiring camera or pose annotations. Our code and project page are available at https://lingtengqiu.github.io/LHM++/.
Figure 1: 3D Avatar Reconstruction and Animation Results of our LHM++. Given a set of $N\geq 1$ images of a human subject, without requiring camera parameters or human pose annotations, our method can reconstruct a high-fidelity, animatable 3D human avatar in seconds.
@@ 1 Introduction
In this work, we propose LHM++, a novel feed-forward framework for fast and high-fidelity 3D human reconstruction from one or a few images without requiring the camera and human poses. To achieve this, we design an efficient Encoder-Decoder Point-Image Transformer (PIT) Framework that hierarchically fuses 3D geometric features with multi-image cues. The framework is built upon Point-Image Transformer blocks (PIT-blocks), which enable interaction between geometric and image tokens via attention fusion while maintaining scalability through spatial hierarchy.
@@
The resulting 3D geometry tokens are decoded to predict Gaussian splatting parameters, enabling high-quality rendering and animation. To further improve visual fidelity, we introduce a lightweight 3D-aware neural animation renderer, built on a DPT-head architecture, that refines the rendering quality of reconstructed avatars in real time. For robustness and generalization, we train our model on large-scale real-world human video datasets spanning diverse clothing styles, body shapes, and viewing conditions. In summary, our contributions are:
  • We introduce LHM++, an efficient feed-forward model, capable of reconstructing high-quality and animatable 3D human avatars in seconds from one or a few casually captured images, without requiring either camera poses or human pose annotations.
@@
  • Extensive experiments on both synthetic and real-world data demonstrate that LHM++ unifies single- and multi-image 3D human reconstruction, with superior generalization and visual quality.
@@ 2 Related Work
Human Reconstruction from A Single Image
@@
Inspired by the success of large reconstruction models [13, 46], emerging solutions aim to enable direct generalizable reconstruction through feed-forward networks which significantly accelerate the inference time. Human-LRM [57] employs a feed-forward model to decode the triplane NeRF representation, then followed by a conditional diffusion-based novel views generation and reconstruction. IDOL [74] introduces a UV-Alignment transformer model to decode Gaussian attribute maps in a structured 2D UV space. LHM [37] leverages a Body-Head multimodal transformer architecture produces animatable 3D avatars with the face identity preservation and fine detail recovery. While these single-view methods often face challenges with occlusions and invisible regions, frequently resulting in geometrically implausible results or blurred textures.
Recently, HumanRAM [67] adapts human reconstruction to novel view and pose synthesis using LVSM with static pose sparse view inputs, achieving photorealistic results but suffering from slow transformer-based rendering. The concurrent work, GIGA [75], employs UV map representations for generalizable reconstruction, but it relies on multi-view captures of the same action, complex camera setups, and motion calibration.
Human Reconstruction from Monocular Videos
@@
Table 1: Comparison with state-of-the-art 3D human reconstruction methods. $F$ stands for Feed-forward, PF for Pose-free, and AM for Animatable.
Method # Image $F$ PF AM Runtime
CAR [25] 11 5 Minutes
IDOL [74] 11 Seconds
AniGS [38] 11 15 Minutes
LHM [37] 11 Seconds
Vid2Avatar [10] >100>100 1-2 Days
Hugs [22] >80>80 30 Minutes
Canonicalfusion [44] >1>1 11 Minutes
GPS-Gaussian [72] 22 Seconds
3DGS-Avatar [34] >20>20 0.5 Hours
InstantAvatar [17] >20>20 Minutes
GaussianAvatar [14] >20>20 0.5-6 Hours
ExAvatar [29] >20>20 1.5-5 Hours
@@
Vid2Avatar-Pro [11] >100>100 Hours
Giga [75] $1\sim 4$ Seconds
FRESA [52] $1\sim 4$ Seconds
@@
Another series of works [17, 29, 14, 11, 18, 35, 66, 68, 65, 22] take it further by incorporating a 3D parametric human model into the optimization process, and thus can drive the human reconstruction in novel poses without any post-processing. Despite impressive visual fidelity, they often require dozens of minutes and dozens of views for a good optimization, which limits their practical usage in real-world scenarios.
Unconstrained collection is ideal input for a practical application. However, existing methods [62, 64] share a similar pipeline that uses a view generative model and score distillation sampling [33] for shape optimization. As a result, they are costly for offline training and impractical for online reconstruction.
Feed-Forward Scene Reconstruction
Recent years have witnessed a paradigm shift in geometric 3D vision, driven by the emergence of methods that eliminate traditional dependencies on camera calibration and multi-stage pipelines. At the forefront of this revolution lies the DUSt3R [53] framework, which reimagines 3D reconstruction as a direct regression problem from image pairs to 3D pointmaps. By discarding the need for intrinsic camera parameters, extrinsic pose estimation, or even known correspondence relationships, DUSt3R and its successors [63, 47, 28, 50] have democratized 3D vision, enabling rapid reconstruction across diverse scenarios while achieving state-of-the-art performance in depth estimation, relative pose recovery, and scene understanding. However, general feed-forward reconstruction methods assume that images are captured from a static scene [51, 23], while our LHM++ can accept human images with different camera and human poses as input and produce an animatable 3D avatar. The concurrent work, FastVGGT [43], employs a token merge strategy to improve the efficiency of image token global attention.
Figure 2: Overview of the proposed LHM++. In 2D space, we extract image tokens $\mathbf{T}_{\text{2D}}$ from input RGB images by DINOv2. In 3D space, geometric tokens $\mathbf{T}_{\text{3D}}$ are derived from SMPL-X anchor points via an MLP. Next, we design an Encoder-Decoder Point-Image Transformer (PIT) to hierarchically fuse 3D and 2D tokens, where the downsampled 3D tokens interact with 2D tokens via multi-modal attention in each layer. The final 3D tokens are decoded to predict 3D Gaussian parameters, followed by a light-weight DPT head for photorealistic animation.
@@ 3.1 Overview
Given a set of $N\geq 1$ RGB images $I^{1},\dots,I^{N}$ of a human subject, without known camera parameters or human pose annotations, our goal is to reconstruct a high-fidelity, animatable 3D human avatar $\chi$ in seconds.
@@
$\mathbf{\chi}\{\mathbf{p},\mathbf{r},\mathbf{f},\rho,\bm{\sigma\}}=\text{LHM++}(P\mid I^{1},\dots,I^{N}).$ (1)
@@
To mitigate this issue, we explore strategies to reduce the number of geometric point tokens involved in attention. However, we empirically observe that simply reducing the number of point tokens significantly degrades reconstruction performance. To address this trade-off, we propose an efficient Encoder-Decoder Point-Image Transformer Framework to fuse image features with geometric point features, as illustrated in Fig. 2, which maintains reconstruction quality while reducing the attention footprint. Additionally, with an increase in input images, image tokens dominate the attention computation. Inspired by ToME [5, 6], we merge the image tokens in a fixed ratio $r$ based on the key similarity map calculated from frame-wise attention before global attention, then unmerge them after global attention. Benefiting from these designs, our model significantly reduces the memory consumption and inference latency and achieves stable performance gains compared to a straightforward extension of LHM.
The final geometric point features output from the decoder are utilized to regress 3D Gaussian parameters using lightweight multi-layer perceptron (MLP) heads. Subsequently, Linear Blend Skinning (LBS) is employed to animate the canonical avatar into the target pose. To enhance the quality of the animation rendering results, we develop a lightweight, real-time neural rendering network that improves rendering outcomes for novel views and poses based on 3D-aware human animation features.
@@ 3.2 Encoder-Decoder Point-Image Transformer Framework
We begin by projecting the SMPL-X anchor points in canonical space into a set of geometric tokens and encoding the input images into image tokens, as described in Sec. 3.3. The encoder-decoder network consists of $N_{\text{layer}}$ PIT blocks. In the first $\lfloor N_{\text{layer}}/2\rfloor$ encoder blocks, we progressively reduce the spatial resolution of the geometric tokens using Grid Pooling [58]. At each layer, the downsampled point tokens perform the attention operation with the image tokens, enabling compact geometric representations enriched with multi-image visual cues.
@@ 3.3 Geometric Point and Images Tokenization
To incorporate human body priors, we initialize a set of 3D query points $\mathbf{X}=\left\{\mathbf{x}_{i}\right\}_{i=1}^{N_{\text{points}}}\subset\mathbb{R}^{3}$ by uniformly sampling from the mesh of a canonical SMPL-X pose. Following the design of Point Transformer v3 (PTv3) [59], we first serialize these points into a structured sequence and then project them into a higher-dimensional feature space using an MLP. Formally, this process is expressed as:
@@
$\mathbf{T}_{\text{I}}=\mathcal{E}_{\text{Img}}(I)\in\mathbb{R}^{N_{\text{I}}\times C},$ (3)
@@ 3.4 Point-Image Transformer Block
Given the image token sequence $\mathbf{T}_{\text{2D}}=\{\mathbf{T}^{1}_{\text{I}},\dots,\mathbf{T}^{N}_{\text{I}}\}\in\mathbb{R}^{N\times N_{\text{I}}\times C}$, we apply self-attention independently to the tokens of each image. This updates the features within each frame based on its own image tokens:
$\mathbf{T}_{\text{2D}},\mathbf{K}_{\text{2D}}=\text{Self-Attention}(\mathbf{T}_{\text{2D}}).$ (5)
where $\mathbf{K}_{\text{2D}}\in\mathbb{R}^{NN_{\text{I}}\times C_{\text{attn}}}$ represents the key features with $C_{\text{attn}}$ channels from QKV in self-attention block, which are used to compute the similarity map of the image tokens.
@@
$\mathbf{\bar{T}}_{\text{3D}},\mathbf{\bar{T}}_{\text{2D}}$ $=\text{MLP}_{\text{proj}}(\mathbf{T}_{\text{3D}}),\text{Merge}(\text{Flatten}(\mathbf{T}_{\text{2D}}),\mathbf{K}_{\text{2D}}),$ (6)
$\mathbf{\bar{T}}_{\text{3D}},\mathbf{\bar{T}}_{\text{2D}}$ $=\text{MM-T}(\mathbf{\bar{T}}_{\text{3D}},\mathbf{\bar{T}}_{\text{2D}}\mid\mathbf{T}_{\text{cls}}),$
$\mathbf{T}_{\text{3D}},\mathbf{T}_{\text{2D}}$ $=\text{MLP}_{\text{uproj}}(\mathbf{\bar{T}}_{\text{3D}}),\text{UnFlatten}(\text{UnMerge}(\mathbf{\bar{T}}_{\text{2D}},\mathbf{K}_{\text{2D}}),N).$
Specifically, $\text{MLP}_{\text{proj}}$ maps point tokens $\mathbf{T}_{\text{3D}}\in\mathbb{R}^{\bar{N}_{\text{points}}\times C_{\text{point}}}$ to $\mathbf{\bar{T}}_{\text{3D}}\in\mathbb{R}^{\bar{N}_{\text{points}}\times C}$, and $\text{MLP}_{\text{uproj}}$ performs the inverse mapping. The Merge and UnMerge operations are adapted from  [6], with a merge ratio $r=0.5$, reducing the number of image tokens from $NN_{\text{I}}$ to $NN_{\text{I}}/r$. This design reduces the time complexity of global attention from $\mathcal{O}((N_{\text{points}}+N)^{2})$ to $\mathcal{O}((\bar{N}_{\text{points}}+N/r)^{2})$.
@@
3.6 3D-Aware Human Animation Rendering
The final output of our model is to drive the reconstructed avatar into view space using the provided human motion control signal, specifically the SMPL-X parameters. Since our Gaussian model is deformed based on SMPL-X, it struggles to accurately represent loose clothing, such as skirts. To improve the rendering quality, we develop a lightweight, 3D-aware DPT head for generating final animation outcomes, inspired by LVSM [19] and HumanRAM [67]. Unlike HumanRAM, which requires processing through transformers with multi-view images for each frame, making it computationally expensive. In contrast, our approach allows for real-time generation.
The neural refinement process can be divided into two parts:
Human Animation Feature Rasterization
As we construct the canonical avatar in canonical space, we can leverage 3D priors to provide fine-grained control information. Specifically, given a reconstructed avatar obtained from Eq.  7 and multi-scale geometric tokens $\mathbf{T}_{\text{3D}}$ from Eq. 6, we employ Gaussian Rasterization [35], denoted as $\text{F}_{\text{GS}}$, to render high-dimensional geometric features into view space, resulting in $\mathbf{I}_{\text{feat}}\in\mathbb{R}^{H\times W\times C_{\textbf{point}}}$. Notably, as no 3D points are available to model the background, we utilize learnable background parameters $\mathbf{F}_{\text{back}}$, to represent background colors. The entire process can be formulated as follows:
$\mathbf{I}^{l}_{\text{feat}}=\text{F}_{\text{GS}}(\mathbf{T}^{l}_{\text{3D}},\mathbf{r},\mathbf{p},\rho,\bm{\sigma};\mathbf{F}_{\text{back}}),$ (8)
where $\mathbf{T}^{l}_{\text{3D}}$ represents the geometry features extracted from the $l$-th decoder layer of the PI Transformer framework.
3D-Aware Neural Animation Rendering
After obtaining the multi-scale 3D-aware human animation feature map $\mathbf{I}^{l}_{\text{feat}}$, we employ a lightweight neural network, as proposed in [39], to fuse these multi-scale animation features for novel view and pose rendering:
$\mathbf{T}^{l}_{\text{feat}}$ $=\text{Patchify}(\mathbf{I}^{l}_{\text{feat}}),$ (9)
$\hat{I}_{\text{NR}},\hat{M}_{\text{NR}}$ $=\text{DPT-head}(\mathbf{T}^{0}_{\text{feat}},\ldots,\mathbf{T}^{{N_{\text{layer}}/2}}_{\text{feat}}),$
where Patchify refers to the operation that converts the 3D-aware animation features into image tokens using $8\times 8$ patch windows, $\hat{I}_{\text{NR}}$ and $\hat{M}_{\text{NR}}$ respectively representing the final predicted RGB image and mask results.